OpenStudy (anonymous):

determine the density of a liquid from the following data . mass of the graduated cynlinder:33.79g mass of the cylinder and liquid:40.14g volume of liquid:13.3ml plz can your help me plz :)

OpenStudy (anonymous):

D= m/v D = (40.14-33.79)/13.3

OpenStudy (anonymous):

You have to subtract out the mass of the cylinder to find the mass of your liquid .

OpenStudy (anonymous):

thnx 1dlove :)